# Adversarial bandit optimization for approximately linear functions

### v3

@@ 1 Introduction
For convex loss functions, the bandit optimization has been extensively studied (see, e.g.,[2, 9, 15]). $\mathcal{O}(d^{1/3}T^{3/4})$ regret bounds are shown by Flaxman et al. [10]. Lattimore shows an information-theoretic regret bound $\widetilde{O}(d^{2.5}\sqrt{T})$ for convex loss functions [14]. For linear loss functions, Abernethy et al. propose SCRIBLE algorithm and give an expected regret bound $\mathcal{O}(d\sqrt{T\ln T})$ [2], achieving optimal dependence on $T$ [8]. Lee et al. propose SCRIBLE with lifting and increasing learning rates and show a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$ [15].
Recently, non-convex functions are also getting popular in this literature. For example, Agarwal et al. show a regret bound $\mathcal{O}(poly(d)T^{2/3})$ for smooth and bounded non-convex functions [4]. Ghai et al. propose algorithms with regret bounds $\mathcal{O}(poly(d)T^{2/3})$ under the assumption that non-convex functions are reparametlized as some convex functions [12].
@@ 2 Related Work
Bandit linear optimization was first proposed by Awerbuch & Kleinberg [6], who achieved a regret bound of $\mathcal{O}(d^{3/5}T^{2/3})$ against oblivious adversaries. Later, McMahan & Blum established a regret bound of $\mathcal{O}(dT^{3/4})$ when facing adaptive adversaries [16]. A foundational approach in bandit optimization problems involves gradient-based smoothing techniques. Abernethy et al. presented pioneering work in this area and achieved an expected regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ when dealing with oblivious adversaries [3]. Bartlett et al. proposed a high-probability regret bound of $\mathcal{O}(d^{2/3}\sqrt{T\ln dT})$ under a special condition [7]. Subsequently, Lee et al. presented a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$ for both oblivious and adaptive adversaries [15].
Unlike bandit convex optimization problems, which have been extensively explored and analyzed, bandit non-convex optimization problems introduce unique challenges due to the complexity of exploring and exploiting in a non-convex area. Gao et al. considered both non-convex losses and non-stationary data and established a regret bound of $O(\sqrt{T+poly(T)}$ [11]. Yang et al. achieved a regret bound of $O(\sqrt{T\log T})$ for non-convex loss functions [19]. However, they both require the loss functions to have smoothness properties, and our loss functions are neither convex nor smooth.
@@
    Revised Generality of Problem Setting: Like the SCRIBLE algorithm, our approach is more general, treating bandit linear optimization as a special case within a broader problem framework.
@@ 3 Preliminaries
This section introduces some necessary notations and defines $\epsilon$-approximately linear function. Then we give our problem setting.
@@ 4.1 SCRIBLE with lifting
We simplify SCRIBLE with lifting and increasing learning rates [15]. The algorithm performs SCRIBLE [2] over a lifted decision set, using a $\nu$-normal barrier $\mathcal{R}$ defined over the $con(\mathcal{K})$ (which always exists) as the regularizer to generate the sequence $x^{\prime}_{1},...,x^{\prime}_{t}$. It set $y^{\prime}_{t}=x^{\prime}_{t}+\mathbf{A}_{t}\mu_{t}$, where $\mathbf{A}_{t}=[\nabla^{2}\mathcal{R}(x^{\prime}_{t})]^{-\frac{1}{2}}$ and $\mu_{t}$ is uniformly sampled at random from $\mathbb{S}_{1}^{d+1}\cap(\mathbf{A}_{t}e_{d+1})^{\perp}$. Since $\mu_{t}$ is orthogonal to $\mathbf{A}_{t}e_{d+1}$, the last coordinate of $\mathbf{A}_{t}\mu_{t}$ is zero, ensuring that $y^{\prime}_{t}=(y_{t},1)$ remains within $\mathcal{K^{\prime}}$. The actual point played is still $y_{t}$. After playing $y_{t}$ and observing $f_{t}(y_{t})(=\theta_{t}^{\top}y_{t}+\sigma_{t}(y_{t}))$, it constructs the loss estimator the same way as SCRIBLE [2]: $g_{t}=df_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$. Furthermore, it adopts the same update method as FTRL algorithm [13]: $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$.
Compared with SCRIBLE with lifting and increasing learning rates [15], which focuses on bandit linear optimization, we do not use the increasing learning rates part, but retain the lifting. This preserves its advantages; for instance, the $\nu$-normal barrier $\mathcal{R}$ always exists on $con(\mathcal{K})$ and we can leverage the properties of the $\nu$-normal barrier. When the loss takes the form $f_{t}(y_{t})=\theta_{t}^{\top}y_{t}$ (i.e., when $\epsilon=0$), the analysis by Lee et al. shows that the first $d$ coordinates of $g_{t}$ are indeed an unbiased estimator of $\theta_{t}$ [15]. However, when $\epsilon\neq 0$ since $g_{t}$ will always be influenced by $\sigma_{t}(y_{t})$, $g_{t}$ is no longer an unbiased estimator of $\theta_{t}$.
@@ 4.2 Analysis
where $h\in\mathcal{K^{\prime}}$, $\theta^{\prime}=(\theta,0)$. Rather than following the decomposition in Lee et al. [15] , divides the regret mainly as:
@@ 5 Proof
This section introduces the preliminary of $\nu$-normal barrier, presents several essential lemmas, and gives the proofs of main theorems.
@@ 5.1 $\nu$-normal barrier
With the help of Equation 8 and Lemma 4, it is easy to apply the triangle inequality to derive $\|x_{t}^{\prime}-h\|_{x_{t}^{\prime}}\leq\|x_{t}^{\prime}\|_{x_{t}^{\prime}}+\|h\|_{x_{t}^{\prime}}\leq\sqrt{\nu}+2\nu$, where $h\in\mathcal{K^{\prime}}$..
@@ 5.2 Useful lemmas
This next lemma is based on Lemma B.9. [15], but due to the differences in the loss functions, $g_{t,i}$ is no longer an unbiased estimate of $\theta_{t,i}$, for $i\in[d]$. Lee et al. state that $\mathbb{E}_{t}[l_{t,i}]=\theta_{t,i}$, for $i\in[d]$ [15], where $l_{t}=d(\theta_{t},0)(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}$. We directly apply it to our analysis.
@@
Recall $\epsilon$-approximately linear function can be write as: $f(x)=\theta^{\top}x+\sigma(x)$. Thus, the regret of SCRIBLE with lifting algorithm
@@
Since $\sigma_{t}$ is chosen after knowing the player’s action, it can cause as large a perturbation as possible. We using $\|\sigma_{t}(x)\|\leq\epsilon$ to bound $\sum_{t=1}^{T}\mathbb{E}[\sigma_{t}(y_{t})-\sum_{t=1}^{T}\sigma_{t}(x^{*})]\leq 2T\epsilon$ and combination of everything, we get
@@ 5.5 Application to black-box optimization
Let $\hat{x}$ be the output of algorithm 1, then from Theorem 1. we can easily prove and ensure $f(\hat{x})-\min_{x\in\mathcal{K}}f(x)\leq\frac{4d\sqrt{2\nu\ln T}}{\sqrt{T}}+\frac{GD}{T^{2}}+d\epsilon(2\nu+\sqrt{\nu})+2\epsilon$. Additionally, we provide a lower bound $2\epsilon$ for this problem (see Lemma 10). We can see that the difference between the lower bound and the upper bound is only $d\epsilon(2\nu+\sqrt{\nu})$ as $T$ approaches infinity. This suggests the potential existence of "easier" settings between the adversarial environment and the standard stochastic environment, where better algorithms might be found. It also motivates us to explore these settings further.

### v4

@@ 1 Introduction
For convex loss functions, the bandit optimization has been extensively studied (see, e.g.,[2, 9, 17]). $\mathcal{O}(d^{1/3}T^{3/4})$ regret bounds are shown by Flaxman et al. [10]. Lattimore [16] shows an information-theoretic regret bound $\widetilde{O}(d^{2.5}\sqrt{T})$ for convex loss functions. For linear loss functions, Abernethy et al. [2] propose the SCRiBLe algorithm and give an expected regret bound $\mathcal{O}(d\sqrt{T\ln T})$, achieving optimal dependence on $T$ [8]. Lee et al. [17] propose SCRiBLe with lifting and increasing learning rates and show a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$.
Recently, non-convex functions are also getting popular in this literature. For example, Agarwal et al. [4] show a regret bound $\mathcal{O}(poly(d)T^{2/3})$ for smooth and bounded non-convex functions. Ghai et al. [12] propose algorithms with regret bounds $\mathcal{O}(poly(d)T^{2/3})$ under the assumption that non-convex functions are reparameterized as some convex functions.
@@ 2 Related Work
Bandit linear optimization was first proposed by Awerbuch & Kleinberg [6], who achieved a regret bound of $\mathcal{O}(d^{3/5}T^{2/3})$ against oblivious adversaries. Later, McMahan & Blum [18] established a regret bound of $\mathcal{O}(dT^{3/4})$ when facing adaptive adversaries. A foundational approach in bandit optimization problems involves gradient-based smoothing techniques. Flaxman et al. [10] proposed a method for constructing an unbiased estimator of the loss function’s gradient in the bandit setting. Abernethy et al. [3] introduced the SCRiBLe algorithm and achieved an expected regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ against oblivious adversaries. Bartlett et al. [7] proposed a high-probability regret bound of $\mathcal{O}(d^{2/3}\sqrt{T\ln dT})$ under a special condition. Subsequently, Lee et al. [17] presented a high-probability regret bound $\widetilde{O}(d^{2}\sqrt{T})$ for both oblivious and adaptive adversaries. In recent work, Ito and Takemura [14, 15] proposed a bandit linear algorithm that adapts to both stochastic and adversarial environments, achieving a regret bound of $\mathcal{O}(d\sqrt{T\ln T})$ in the adversarial setting. Rodemann et al. [21] established a connection between bandit problems and Bayesian black-box optimization, offering theoretical foundations for regret bounds across both domains.
Unlike bandit convex optimization problems, which have been extensively explored and analyzed, bandit non-convex optimization problems introduce unique challenges due to the complexity of exploring and exploiting in a non-convex area. Gao et al. [11] considered both non-convex losses and non-stationary data and established a regret bound of $O(\sqrt{T+poly(T)}$. Yang et al. [22] achieved a regret bound of $O(\sqrt{T\log T})$ for non-convex loss functions. However, they both require the loss functions to have smoothness properties, and our loss functions are neither convex nor smooth.
@@
    Revised Generality of Problem Setting: Like the SCRiBLe algorithm [3], our approach is more general, treating bandit linear optimization as a special case within a broader problem framework.
@@ 3 Preliminaries
This section introduces some necessary notations and defines an $\epsilon$-approximately linear function. Then we give our problem setting.
@@ 4.1 SCRiBLe with lifting
We simplify SCRiBLe with lifting and increasing learning rates [17]. The algorithm performs original SCRiBLe [2] over a lifted decision set, using a $\nu$-normal barrier $\mathcal{R}$ defined over the $con(\mathcal{K})$ (which always exists) as the regularizer to generate the sequence $x^{\prime}_{1},...,x^{\prime}_{t}$. It set $y^{\prime}_{t}=x^{\prime}_{t}+\mathbf{A}_{t}\mu_{t}$, where $\mathbf{A}_{t}=[\nabla^{2}\mathcal{R}(x^{\prime}_{t})]^{-\frac{1}{2}}$ and $\mu_{t}$ is uniformly sampled at random from $\mathbb{S}_{1}^{d+1}\cap(\mathbf{A}_{t}e_{d+1})^{\perp}$. Since $\mu_{t}$ is orthogonal to $\mathbf{A}_{t}e_{d+1}$, the last coordinate of $\mathbf{A}_{t}\mu_{t}$ is zero, ensuring that $y^{\prime}_{t}=(y_{t},1)$ remains within $\mathcal{K^{\prime}}$. The actual point played is still $y_{t}$. After playing $y_{t}$ and observing $f_{t}(y_{t})(=\theta_{t}^{\top}y_{t}+\sigma_{t}(y_{t}))$, it constructs the loss estimator the same way as SCRiBLe algorithm [2]: $g_{t}=df_{t}(y_{t})\mathbf{A}_{t}^{-1}\mu_{t}$. Furthermore, it adopts the same update method as FTRL algorithm [13]: $x^{\prime}_{t+1}=\arg\min\limits_{x^{\prime}\in\mathcal{K^{\prime}}}{\eta\sum_{\tau=1}^{t}g_{\tau}^{\top}x^{\prime}+\mathcal{R}(x^{\prime})}$.
Compared with SCRiBLe with lifting and increasing learning rates [17], which focuses on bandit linear optimization, we do not use the increasing learning rates part, but retain the lifting. This preserves its advantages; for instance, the $\nu$-normal barrier $\mathcal{R}$ always exists on $con(\mathcal{K})$ and we can leverage the properties of the $\nu$-normal barrier. When the loss takes the form $f_{t}(y_{t})=\theta_{t}^{\top}y_{t}$ (i.e., when $\epsilon=0$), the analysis by Lee et al. [17] shows that the first $d$ coordinates of $g_{t}$ are indeed an unbiased estimator of $\theta_{t}$. However, when $\epsilon\neq 0$ since $g_{t}$ will always be influenced by $\sigma_{t}(y_{t})$, $g_{t}$ is no longer an unbiased estimator of $\theta_{t}$.
@@ 4.2 Analysis
where $h\in\mathcal{K^{\prime}}$, $\theta^{\prime}=(\theta,0)$. This differs from the decomposition proposed by Lee et al. [17], which divides the regret as follows:
@@ 5 Proof
This section introduces preliminaries on the $\nu$-normal barrier, presents several essential lemmas and provides proofs of the main theorems.
@@ 5.1 $\nu$-normal barrier
With the help of Equation 8 and Lemma 4, it is easy to apply the triangle inequality to derive $\|x_{t}^{\prime}-h\|_{x_{t}^{\prime}}\leq\|x_{t}^{\prime}\|_{x_{t}^{\prime}}+\|h\|_{x_{t}^{\prime}}\leq\sqrt{\nu}+2\nu$, where $h\in\mathcal{K^{\prime}}$.
@@ 5.2 Useful lemmas
This next lemma is based on Lemma B.9. [17], but due to the differences in the loss functions, $g_{t,i}$ is no longer an unbiased estimate of $\theta_{t,i}$, for $i\in[d]$. Lee et al. state that $\mathbb{E}_{t}[l_{t,i}]=\theta_{t,i}$, for $i\in[d]$, where $l_{t}=d(\theta_{t},0)(x_{t}^{\prime}+\mathbf{A}_{t}\mu_{t})\mathbf{A}_{t}^{-1}\mu_{t}$ [17]. We directly apply it to our analysis.
@@
Recall an $\epsilon$-approximately linear function can be written as: $f(x)=\theta^{\top}x+\sigma(x)$. Thus, the regret of SCRiBLe with lifting algorithm
@@
Since $\sigma_{t}$ is chosen after knowing the player’s action, it can cause as large a perturbation as possible. We using $\|\sigma_{t}(x)\|\leq\epsilon$ to bound $\sum_{t=1}^{T}\mathbb{E}[\sigma_{t}(y_{t})-\sum_{t=1}^{T}\sigma_{t}(x^{*})]\leq 2T\epsilon$ and combination of Lemma 8, we get
@@ 5.5 Application to black-box optimization
Let $\hat{x}$ be the output of Algorithm 1. Then, it follows from Theorem 1 that $f(\hat{x})-\min_{x\in\mathcal{K}}f(x)\leq\frac{4d\sqrt{2\nu\ln T}}{\sqrt{T}}+\frac{GD}{T^{2}}+d\epsilon(2\nu+\sqrt{\nu})+2\epsilon$. Additionally, we provide a lower bound $2\epsilon$ for this problem (see Lemma 10). We can see that the difference between the lower bound and the upper bound is only $d\epsilon(2\nu+\sqrt{\nu})$ as $T$ approaches infinity. This suggests the potential existence of "easier" settings between the adversarial environment and the standard stochastic environment, where better algorithms might be found. It also motivates us to explore these settings further.
@@ 7 Experiments
8 Conclusion
In this work, we investigate the bandit optimization problem involving a class of non-convex and non-smooth loss functions, where in each trial the loss function is the sum of a linear function and a small but arbitrary perturbation chosen after observing the player’s choice. We introduce a novel approach for analyzing high-probability regret and derive both upper and lower bounds for the problem. For future work, we aim to study the bandit optimization problem for a broader class of non-convex and non-smooth loss functions, such as those formed by the sum of a convex function and a perturbation.